# Enhancing Metabolic Syndrome Prediction with Hybrid Data Balancing and Counterfactuals

### v1

@@ Abstract
Metabolic Syndrome (MetS) is a cluster of interrelated risk factors that significantly increases the risk of cardiovascular diseases and type 2 diabetes. Despite its global prevalence, accurate prediction of MetS remains challenging due to issues such as class imbalance, data scarcity, and methodological inconsistencies in existing studies. In this paper, we address these challenges by systematically evaluating and optimizing machine learning (ML) models for MetS prediction, leveraging advanced data balancing techniques and counterfactual analysis. Multiple ML models, including XGBoost, Random Forest, TabNet, etc., were trained and compared under various data balancing techniques such as random oversampling (ROS), SMOTE, ADASYN, and CTGAN. Additionally, we introduce MetaBoost [^3] The source code is available at: github.com/asusanyamshah/MetaBoost, a novel hybrid framework that integrates SMOTE, ADASYN, and CTGAN, optimizing synthetic data generation through weighted averaging and iterative weight tuning to enhance the model’s performance (achieving a 1.14% accuracy improvement over individual balancing techniques). A comprehensive counterfactual analysis is conducted to quantify feature-level changes required to shift individuals from high-risk to low-risk categories. The results indicate that blood glucose (50.3%) and triglycerides (46.7%) were the most frequently modified features, highlighting their clinical significance in MetS risk reduction. Additionally, probabilistic analysis shows elevated blood glucose (85.5% likelihood) and triglycerides (74.9% posterior probability) as the strongest predictors. This study not only advances the methodological rigor of MetS prediction but also provides actionable insights for clinicians and researchers, highlighting the potential of ML in mitigating the public health burden of metabolic syndrome.
@@ I Introduction
Challenges in healthcare datasets, such as missing or imbalanced data and small sample sizes require the development of robust predictive models for accurate inference [5, 6, 7, 8]. To address these challenges, techniques such as random oversampling (ROS), Synthetic Minority Oversampling Technique (SMOTE) [9], Adaptive Synthetic Sampling (ADASYN), and generative models like Conditional Tabular Generative Adversarial Networks (CTGAN) have been employed for data balancing [10, 11]. However, the comparative efficacy of these methods, particularly in combination, remains underexplored in the context of MetS prediction.
Recent advances in ML offer the transformative potential for improving MetS prediction by integrating heterogeneous clinical, demographic, and laboratory data into holistic risk assessments [12]. Despite this promise, existing studies often face limitations such as class imbalance, data scarcity, and inconsistent preprocessing protocols, which can lead to inflated performance metrics and reduced generalizability [13]. While ML models like XGBoost, Random Forest, and deep learning architectures have demonstrated promise in metabolic health analytics, rigorous comparative analyses under standardized preprocessing and evaluation frameworks are lacking.
@@
To bridge these gaps, we investigate the following research question: How can hybrid machine learning approaches, leveraging advanced data balancing techniques and counterfactual analysis, enhance the early detection, risk stratification, and clinical interpretability of MetS prediction models? Motivated by these gaps, this study aims to systematically evaluate and optimize ML models for MetS prediction, leveraging a comprehensive preprocessing pipeline and advanced data balancing techniques. Specifically, we address class imbalance through a combination of different methods for oversampling, such as SMOTE, ADASYN, and CTGAN, exploring both individual and hybrid approaches to synthetic data generation. To this end, we introduce MetaBoost, a novel hybrid framework that integrates these techniques through weighted averaging and iterative weight tuning to optimize synthetic data generation and improve model robustness. Furthermore, we conduct a detailed counterfactual analysis to elucidate the feature-level changes required to shift individuals from high-risk to low-risk categories, providing actionable insights for clinical interventions.
@@ II-A Dataset
The dataset used for this experiment comprises demographic, clinical, and laboratory measurements for 2,402 individuals, with 15 features, including the response variable indicating the presence or absence of metabolic syndrome. A risk analysis was conducted to assess the prevalence of metabolic syndrome in the dataset and to compute likelihood and posterior probabilities for specific feature combinations.
@@ II-B Data Preprocessing
The preprocessing pipeline involved several critical steps to ensure the dataset was suitable for analysis. After removing columns with low relevance, categorical variables, such as “sex” and “race” were encoded numerically. For “sex”, Male was encoded as 0 and Female as 1. For “race”, the categories were mapped as follows: White (0), Asian (1), Black (2), Mexican American (3), Hispanic (4), and Other (5).
Missing values in the columns “Income”, “WaistCirc”, and “BMI” were imputed using mean imputation. The dataset was then split into training and testing sets, with the testing set comprising 30% of the total data. It was ensured that the test set contained an equal number of instances for both classes of the response variable (presence or absence of metabolic syndrome) to avoid bias in evaluation.
@@ II-D Analysis of Oversampling Strategies
Three distinct strategies were employed to address class imbalance. The first included applying random oversampling to the entire dataset before splitting it into training and testing sets. We understand that the first method has flaws such as the risk of data leakage and also being evaluated on synthesized data, we kept this analysis for establishing an upper bound of the results. The second approach did not use any oversampling and models were trained on the original imbalanced dataset. Finally, in the third approach, random oversampling was applied only to the training set, while the testing set remained unchanged. For each strategy, the dataset was split such that the testing set constituted 30% of the total data, with equal representation of both classes. Performance metrics (accuracy, precision, recall, and F1 score) were recorded for each model and strategy combination.
@@ II-E Analysis of Class Imbalance Handling Methods
To further mitigate class imbalance, advanced techniques such as Synthetic Minority Oversampling Technique (SMOTE), Adaptive Synthetic Sampling (ADASYN), and Conditional Tabular Generative Adversarial Networks (CTGAN) were employed  [16]. These methods were tested individually and in hybrid combinations. For hybrid approaches, synthetic data points were generated by calculating a weighted average of the nearest pairs from the training dataset. For combinations involving two methods the weights for each method were systematically varied in increments of 0.05, with 20 iterations performed for each combination. For hybrid approaches combining two methods that involved CTGAN, CTGAN was trained for 300 epochs.
@@ II-F Counterfactual Analysis
Counterfactual examples were generated using the Neural Importance Counterfactual Explanations (NICE) [17] algorithm to identify the minimal feature modifications required to shift individuals between the positive class (MetS present) and negative class (MetS absent) [18]. The counterfactual search was formulated as an optimization problem:
$\arg\min_{x^{\prime}}\mathcal{D}(x,x^{\prime})+\lambda\cdot\mathcal{C}(x^{\prime})$ (1)
where $x$ represents the original instance, and $x^{\prime}$ is the counterfactual example. $\mathcal{D}(x,x^{\prime})$ measures the feature-wise distance between instances, computed using the L1 norm. $\mathcal{C}(x^{\prime})$ ensures that $x^{\prime}$ belongs to the opposite class of $x$ (i.e., $f(x)\neq f(x^{\prime})$). $\lambda$ is a regularization parameter balancing proximity and classification change. To quantify feature modifications, a weighted feature-level distance was employed:
$\mathcal{D}_{w}(x,x^{\prime})=\sum_{i=1}^{d}w_{i}|x_{i}-x^{\prime}_{i}|$ (2)
where $w_{i}$ represents the normalized importance of each feature, derived from model-specific feature importance metrics such as SHAP values or Random Forest impurity scores.
@@ III-A1 Prior Probability
Parameter Measure Likelihood %
glucose_risk 11 85.5%
bmi_risk 8 62.4%
triglycerides_risk 14 54.7%
uralbcr_risk 16 20.3%
albuminuria_risk 15 15.9%
@@
Parameter Measure Posterior %
triglycerides_risk 8 74.9%
bmi_risk 2 59.9%
glucose_risk 5 58.7%
uralbcr_risk 10 53.5%
albuminuria_risk 9 51.6%
@@ III-A2 Likelihood Analysis
TABLE III: Performance of models across different random oversampling (ROS) strategies. Values in () show performance without any ROS. The best values are highlighted in bold. Acc: Accuracy, Pre: Precision, Rec: Recall, F1: F1 Score. RF: Random Forest, DT: Decision Tree, LR: Logistic Regression, TNet: TabNet.
@@ III-A3 Posterior Probabilities
The results shown in Table  III are as expected. The accuracy when ROS is performed on the entire dataset is the highest since data leakage is occurring. The accuracy when ROS is not applied is lower than when ROS is applied only on the training set. This shows that ROS is working and corrects the limitations posed by the class imbalance in the data set.
The results demonstrated noticeable variations in model performance across different random over-sampling strategies. Key observations include:
@@ III-B Accuracy Comparison
The XGBoost classifier consistently showed the highest accuracy. However, MLP consistently performed the worst in terms of accuracy. This might be due to the complexity of MLP, which might require more data for MLP to perform well. MLP is also prone to overfitting, which might explain its poor metrics [26].
@@ III-C Precision, Recall, and F1
The XGBoost Classifier consistently achieves the highest precision and F1 score. In contrast, the MLP shows variable precision due to its sensitivity to input feature scaling and tendency to overfit [26]. Notably, TabNet achieves the highest recall both with and without ROS, thanks to its sequential attention mechanism, effective handling of imbalanced datasets, and hybrid approach combining decision trees and neural networks. These features enhance TabNet’s ability to identify true positives and generalize well, while regularization techniques prevent overfitting. Conversely, the MLP exhibits the lowest F1 scores, particularly without ROS, due to its overfitting tendencies and the need for extensive tuning and large datasets [26].
TABLE IV: Comparison of test accuracy, precision, recall, and F1 score of hybrid data balancing approaches with their optimal weights. Data generation was applied on the training dataset and the testing was done only on real data.
@@
III-D Correct Analysis - No Synthetic Data In the Test Set
In the first strategy, the training and testing data were separated after performing ROS on the entire dataset. This approach has flaws since some synthetic training and test data may be generated from the same sources or neighbors. Not performing ROS is also impractical since the data imbalance problem remains.
The final approach was implemented by first manually splitting the training and testing data so that the testing set contains 30% of the data points from the original dataset. Both the classes will be equally represented in the testing set. Then, random over sampling was performed on only the training data. Considering the data from this analysis, XGBoost has the highest accuracy and F1 score. MLP had the lowest accuracy, precision and F1 score. This result can be attributed to its susceptibility to overfitting and complexity, which requires a lot of data for it to have good performance  [26]. Random Forest had the highest precision. MLP had the highest recall.
@@ III-E MetaBoost - Combining Multiple Data Balancing Methods
When multiple balancing methods were combined using the methods described above, the highest F1 score was achieved by a combination of ADASYN and CTGAN with weights of $0.4$ and $0.6$, respectively. This combination yielded an accuracy of $87.13\%$. The higher weight assigned to CTGAN ($0.6$) compared to ADASYN ($0.4$) indicates that CTGAN performed slightly better than ADASYN in this context.
For the combination of SMOTE and CTGAN, the highest F1 score was achieved when both methods were assigned equal weights ($0.5,0.5$). This suggests that CTGAN performed just as well as SMOTE, implying that SMOTE is slightly more effective than ADASYN in this approach. This observation is further supported by the results of combining SMOTE and ADASYN, where SMOTE was assigned a higher weight of $0.75$ compared to ADASYN’s $0.25$.
The combination of all three methods (SMOTE, CTGAN, and ADASYN) with weights $0.05$, $0.55$, and $0.4$, respectively, resulted in an accuracy of $86.88\%$. This indicates that CTGAN still plays a dominant role in the combined approach, while SMOTE and ADASYN contribute to a lesser extent. Interestingly, here ADASYN is assigned a much higher weight than SMOTE. This might be because ADASYN might be providing synthetic samples that complement those generated by CTGAN and SMOTE effectively for this combination of weights. This complementary effect can enhance the overall diversity and quality of the synthetic data, leading to better model performance. There might also be interaction effects between the synthetic samples generated by the different methods that might explain the unpredictable nature of the weights that result in the best performance. This is presented in Table  IV and Fig. 2.
Figure 2: Comparison of accuracies of hybrid data balancing approaches across different weights.
@@ III-G Evaluation of the Counterfactual Explanations
$\mathbb{E}\left[\frac{||x-x^{\prime}||_{0}}{d}\right]=\frac{1}{N}\sum_{j=1}^{N}\frac{\sum_{i=1}^{d}\mathbf{1}(x_{i,j}\neq x^{\prime}_{i,j})}{d}$ (3)
@@
TABLE V: Summary metrics of counterfactual examples. This table quantifies the feature modifications required to shift individuals from high-risk to low-risk MetS categories
@@ III-H Feature-Specific Analysis
As presented in Table  VI the feature-specific analysis shows that metabolic markers were most frequently modified. Blood Glucose (50.3%) and Triglycerides (46.7%) exhibited the highest change rates, followed by Waist Circumference (42.9%) and HDL (33.7%). This pattern aligns with clinical understanding, as these features are crucial metabolic indicators. Notably, demographic features (Sex (0.1%), Race (0.0%)), and socioeconomic factors (Income (1.7%)) showed minimal modifications, suggesting the model’s predictions are robust to these characteristics. The relatively low modification rates for Age (8.9%) and anthropometric measures like BMI (9.6%) indicate that while these features are clinically relevant, they require less adjustment to affect the model’s predictions. This analysis reveals that the model primarily relies on modifiable metabolic factors rather than fixed demographic characteristics for its predictions which proves its usefulness in mitigating risk for metabolic syndrome through recommendations for achievable changes.
TABLE VI: Feature-specific change rates in counterfactual analysis. This table highlights the most frequently modified features in counterfactual examples, with blood glucose (50.3%) and triglycerides (46.7%) being the most altered, reinforcing their critical role in MetS risk reduction.

### v2

@@ Abstract
Metabolic Syndrome (MetS) is a cluster of interrelated risk factors that significantly increases the risk of cardiovascular diseases and type 2 diabetes. Despite its global prevalence, accurate prediction of MetS remains challenging due to issues such as class imbalance, data scarcity, and methodological inconsistencies in existing studies. In this paper, we address these challenges by systematically evaluating and optimizing machine learning (ML) models for MetS prediction, leveraging advanced data balancing techniques and counterfactual analysis. Multiple ML models, including XGBoost, Random Forest, TabNet, etc., were trained and compared under various data balancing techniques such as random oversampling (ROS), SMOTE, ADASYN, and CTGAN. Additionally, we introduce MetaBoost [^3] The source code is available at: github.com/asusanyamshah/MetaBoost, a novel hybrid framework that integrates SMOTE, ADASYN, and CTGAN, optimizing synthetic data generation through weighted averaging and iterative weight tuning to enhance the model’s performance (achieving up to a 1.87% accuracy improvement over individual balancing techniques). A comprehensive counterfactual analysis is conducted to quantify the feature-level changes required to shift individuals from high-risk to low-risk categories. The results indicate that blood glucose (50.3%) and triglycerides (46.7%) were the most frequently modified features, highlighting their clinical significance in MetS risk reduction. Additionally, probabilistic analysis shows elevated blood glucose (85.5% likelihood) and triglycerides (74.9% posterior probability) as the strongest predictors. This study not only advances the methodological rigor of MetS prediction but also provides actionable insights for clinicians and researchers, highlighting the potential of ML in mitigating the public health burden of metabolic syndrome.
@@ I Introduction
Challenges in healthcare datasets, such as missing or imbalanced data and small sample sizes, require the development of robust predictive models for accurate inference [5, 6, 7, 8, 9]. To address these challenges, techniques such as random oversampling (ROS), Synthetic Minority Oversampling Technique (SMOTE) [10], Adaptive Synthetic Sampling (ADASYN) [11], and generative models like Conditional Tabular Generative Adversarial Networks (CTGAN) [12], BIDC2 [13], and AIMEN [14] have been employed for data balancing. During synthetic data generation, it is important not to add too much post-generation filitering as it can cause the distribution of the training set to be different than that of the test set [14]. However, the comparative efficacy of these methods, particularly in combination, remains underexplored in the context of MetS prediction.
Recent advances in ML offer the transformative potential for improving MetS prediction by integrating heterogeneous clinical, demographic, and laboratory data into holistic risk assessments [15, 16]. Existing works on hyperglycemia and hypoglycemia detection demonstrate the potential of leveraging machine learning and simulation-based approaches for personalized predictions and tailored interventions in metabolic health [17, 18, 19]. Existing studies often face limitations such as class imbalance, data scarcity, and inconsistent preprocessing protocols, which can lead to inflated performance metrics and reduced generalizability [20, 21, 22]. While ML models like XGBoost, Random Forest, and deep learning architectures have demonstrated promise in metabolic health analytics, rigorous comparative analyses under standardized preprocessing and evaluation frameworks are lacking.
@@
To bridge these gaps, we investigate the following research question: How can hybrid machine learning approaches, leveraging advanced data balancing techniques and counterfactual analysis, enhance the early detection, risk stratification, and clinical interpretability of MetS prediction models? Motivated by these gaps, this study aims to systematically evaluate and optimize ML models for MetS prediction, leveraging a comprehensive preprocessing pipeline and advanced data balancing techniques. Specifically, we address class imbalance through a combination of different methods for oversampling, such as SMOTE, ADASYN, and CTGAN, exploring both individual and hybrid approaches to synthetic data generation. To this end, we introduce MetaBoost, a novel hybrid framework that integrates these techniques through weighted averaging and iterative weight tuning to optimize synthetic data generation and improve model robustness. Furthermore, we conduct a counterfactual analysis to elucidate the feature-level changes required to shift individuals from high-risk to low-risk categories, providing actionable insights for clinical interventions.
@@ II-A Dataset
The dataset used for this experiment comprises demographic, clinical, and laboratory measurements for 2,401 individuals, with 15 columns that entails 13 features, a sequence ID and the response variable indicating the presence or absence of metabolic syndrome. The 13 features include age, sex, marital status, income, race, waist circumference, BMI (body mass index), albuminuria (presence of albumin, a blood protein, in the urine), urine albumin–creatinine ratio, uric acid, blood glucose, HDL (high-density lipoprotein), and triglycerides. This is a dataset derived from the National Health and Nutrition Examination Survey (NHANES) database. We refer the readers to the original data collection paper for further details about the dataset and the data collection procedure [23]. We conducted a risk analysis to assess the prevalence of metabolic syndrome in the dataset and to compute likelihood and posterior probabilities for specific feature combinations.
@@ II-B Data Preprocessing
The preprocessing pipeline involved several critical steps to ensure the dataset was suitable for analysis. As a considerable number of rows (8.66%) had missing values for marital status, this feature was removed from the input feature set. Categorical variables, such as “sex” and “race” were encoded numerically. For “sex”, Male was encoded as 0 and Female as 1. For “race”, the categories were mapped as follows: White (0), Asian (1), Black (2), Mexican American (3), Hispanic (4), and Other (5).
Missing values in the columns “Income”, “WaistCirc”, and “BMI” were imputed using mean imputation. The dataset was then split into training and testing sets, with the testing set comprising 33% of the total data. It was ensured that the test set contained an equal number of instances for both classes of the response variable (presence or absence of metabolic syndrome) to avoid misinterpretation of the evaluation results.
@@ II-D Analysis of Oversampling Strategies
Two distinct strategies were employed to address class imbalance. The first approach did not use any oversampling, and models were trained on the original imbalanced dataset. In the second approach, random oversampling was applied only to the training set, while the testing set remained unchanged. For each strategy, the dataset was split such that the testing set constituted 33% of the total data, with equal representation of both classes in the test set. Performance metrics (accuracy, precision, recall, and F1 score) were recorded for each model and strategy combination.
@@ II-E Analysis of Class Imbalance Handling Methods
To further mitigate class imbalance, advanced techniques such as Synthetic Minority Oversampling Technique (SMOTE), Adaptive Synthetic Sampling (ADASYN), and Conditional Tabular Generative Adversarial Networks (CTGAN) were employed  [26]. These methods were tested individually and in hybrid combinations. For hybrid approaches, synthetic data points were generated by calculating a weighted average of the nearest pairs from the training dataset. For combinations involving two methods, the weights for each method were systematically varied in increments of 0.05, with 20 iterations performed for each combination. For hybrid approaches combining two methods that involved CTGAN, CTGAN was trained for 300 epochs.
@@ II-F Counterfactual Analysis
Counterfactual examples were generated using the Nearest Instance Counterfactual Explanations (NICE) [27] algorithm to identify the minimal feature modifications required to shift individuals between the positive class (MetS present) and negative class (MetS absent). The counterfactual search was formulated as an optimization problem:
$\vspace{-1mm}\arg\min_{x^{\prime}}\mathcal{D}(x,x^{\prime})+\lambda\cdot\mathcal{C}(x^{\prime})$ (1)
where $x$ represents the original instance, and $x^{\prime}$ is the counterfactual example. $\mathcal{D}(x,x^{\prime})$ measures the feature-wise distance between instances, computed using the L1 norm. $\mathcal{C}(x^{\prime})$ ensures that $x^{\prime}$ belongs to the opposite class of $x$ (i.e., $f(x)\neq f(x^{\prime})$). $\lambda$ is a regularization parameter balancing proximity and classification change.
@@ III-A1 Prior Probability
Parameter Likelihood %
glucose_risk 85.5%
bmi_risk 62.4%
triglycerides_risk 54.7%
uralbcr_risk 20.3%
albuminuria_risk 15.9%
@@
Parameter Posterior %
triglycerides_risk 74.9%
bmi_risk 59.9%
glucose_risk 58.7%
uralbcr_risk 53.5%
albuminuria_risk 51.6%
@@ III-A2 Likelihood Analysis
TABLE III: Performance of models across different random oversampling (ROS) strategies. Values in the first section show performance without any ROS. The best values are highlighted in bold. Acc: Accuracy, Pre: Precision, Rec: Recall, F1: F1 Score. RF: Random Forest, DT: Decision Tree, LR: Logistic Regression, TNet: TabNet.
@@ III-B Accuracy Comparison
The results shown in Table  III are consistent with our hypothesis that training an imbalanced dataset, where no ROS or other balancing methods were performed, would yield lower accuracy and F1 score. The best test accuracy when ROS is not applied (0.843) is lower than the test accuracy when ROS is applied to the training set (0.859).
The results demonstrated noticeable variations in model performance across different random over-sampling strategies. One key observation is that the XGBoost classifier consistently showed the highest accuracy. However, MLP consistently performed the worst in terms of accuracy. This might be due to the complexity of MLP, which might require more data for MLP to perform well. MLP is also prone to overfitting, which might explain its poor metrics [35].
@@ III-C Precision, Recall, and F1
The XGBoost Classifier consistently achieves the highest precision and F1 score. In contrast, the MLP shows variable precision due to its sensitivity to input feature scaling and tendency to overfit [35]. Notably, TabNet achieves the highest recall without ROS, thanks to its sequential attention mechanism, effective handling of imbalanced datasets, and hybrid approach combining decision trees and neural networks. These features enhance TabNet’s ability to identify true positives and generalize well, while regularization techniques prevent overfitting. Conversely, the MLP exhibits the lowest F1 scores, particularly without ROS, due to its overfitting tendencies and the need for extensive tuning and large datasets [35].
TABLE IV: Comparison of test accuracy, precision, recall, and F1 score of hybrid data balancing approaches with their optimal weights. Data generation was applied to the training dataset, and the testing was done only on real data.
@@
III-D Analysis with ROS on the Training Set
The first approach of Table  III, where ROS is not performed, keeps the training data imbalanced, and achieving a good performance on the test set is often impractical in this way. The other approach was implemented by first manually splitting the training and testing data so that the testing set contains 33% of the data points from the original dataset. Both classes would be equally represented in the testing set. Then, random oversampling was performed on only the training data. Considering the data from this analysis, XGBoost has the highest accuracy (0.859) and F1 score (0.849). MLP had the lowest accuracy, precision, and F1 score but the highest recall. On the other hand, Random Forest had the highest precision.
@@ III-E MetaBoost - Combining Multiple Data Balancing Methods
When multiple balancing methods were combined using the methods described above, the highest F1 score was achieved by a combination of ADASYN and CTGAN with weights of $0.4$ and $0.6$, respectively. This combination yielded an accuracy of $87.1\%$ and an F1 score of 0.868. The higher weight assigned to CTGAN ($0.6$) compared to ADASYN ($0.4$) indicates that CTGAN has contributed more than ADASYN to the generated synthetic training data in this context. Moreover, the hybrid data balancing method (F1 score: 0.868) outperforms the best result found with individual balancing methods (F1 score: 0.864).
For the combination of SMOTE and CTGAN, the highest F1 score (0.865) was achieved when both methods were assigned equal weights ($0.5,0.5$). This suggests that CTGAN and SMOTE had equal contributions or priorities in the synthetic training data samples. Finally, with the training dataset combining SMOTE and ADASYN, SMOTE was assigned a higher weight of $0.75$ compared to ADASYN’s $0.25$ to reach the highest F1 score (0.858) of that combination.
The combination of all three methods (SMOTE, CTGAN, and ADASYN) with weights $0.05$, $0.55$, and $0.4$, respectively, resulted in an accuracy of $86.88\%$. This indicates that CTGAN still plays a dominant role in the combined approach, while SMOTE and ADASYN contribute to it to a lesser extent. Interestingly, here ADASYN is assigned a much higher weight than SMOTE. This might be because ADASYN might be providing synthetic samples that complement those generated by CTGAN and SMOTE effectively for this combination of weights. This complementary effect can enhance the overall diversity and quality of the synthetic data, leading to better model performance. There might also be interaction effects between the synthetic samples generated by the different methods that might explain the unpredictable nature of the weights that result in the best performance. This is presented in Table  IV and Fig. 2.
Figure 2: Comparison of accuracies of hybrid data balancing approaches across different weights. Weights: ($\lambda,1-\lambda$).
@@ III-G Evaluation of the Counterfactual Explanations
$\vspace{-2mm}\mathbb{E}\left[\frac{||x-x^{\prime}||_{0}}{d}\right]=\frac{1}{N}\sum_{j=1}^{N}\frac{\sum_{i=1}^{d}\mathbf{1}(x_{i,j}\neq x^{\prime}_{i,j})}{d}$ (2)
@@
TABLE V: Summary metrics of counterfactual examples. This table quantifies the feature modifications required to shift individuals from high-risk to low-risk MetS categories.
@@ III-H Feature-Specific Analysis
As presented in Table  VI, the feature-specific analysis shows that metabolic markers were most frequently modified. Blood Glucose (50.3%) and Triglycerides (46.7%) exhibited the highest change rates, followed by Waist Circumference (42.9%) and HDL (33.7%). This pattern aligns with clinical understanding, as these features are crucial metabolic indicators. Notably, demographic features (Sex (0.1%), Race (0.0%), and socioeconomic factors (Income (1.7%) showed minimal modifications, suggesting the model’s predictions are robust to these characteristics. The relatively low modification rates for Age (8.9%) and anthropometric measures like BMI (9.6%) indicate that while these features are clinically relevant, they require less adjustment to affect the model’s predictions. This analysis reveals that the model primarily relies on modifiable metabolic factors rather than fixed demographic characteristics for its predictions, which proves its usefulness in mitigating risk for metabolic syndrome through recommendations for achievable changes.
TABLE VI: Feature-specific change rates in counterfactual analysis. This table highlights the most frequently modified features in counterfactual examples, with blood glucose and triglycerides being the most altered, reinforcing their critical role in MetS risk reduction.